# SUREL+: Moving from Walks to Sets for Scalable Subgraph-based Graph Representation Learning

### 17 Dec 2023 v2

@@ 3. The framework of SUREL+
This section introduces SUREL+, whose key concept is to offline sample node sets around seed nodes in the graph, which can be joined online as a proxy of query-induced subgraph for representation learning. This approach keeps only distinct nodes in the sampled set for reuse in different queries, effectively addressing memory and computation concerns of node duplication in existing walk-based representations adopted by SUREL (Yin et al., 2022). SUREL+ features a modular design that supports various set samplers and structure encoders, and different set neural encoders can be paired with diverse structural features to compensate for the structure information loss after reducing subgraphs to node sets. Furthermore, SUREL+ introduces a customized sparse data structure SpG and an arithmetic operator SpJoin to store node sets and perform their online joins efficiently. Fig. 2 summarizes and compares SUREL+ and current SGRL models. The following subsections describe these modules in detail.
@@ 3.1. Set Samplers and Structure Encoders
SUREL+ uses set samplers to sample a set of nodes from the neighborhood of each seed node in the graph and calls structure encoders to construct structural features. Both operations are executed offline: the former is primarily for computational benefits, while the latter is to offset the structure loss of node sets reducing from subgraphs (adopted by SEAL (Zhang and Chen, 2018; Zhang et al., 2021)) or walks (adopted by SUREL (Yin et al., 2022)). Conceptually, SUREL+ represents the node-induced subgraph $\mathcal{G}_{u}$ via a combination of (1) a node set $\mathcal{S}_{u}$ comprising unique nodes sampled from the neighborhood of node $u$ and (2) the associated structural features $\mathcal{Z}_{u}$ reflects the position of sampled nodes in $\mathcal{G}_{u}$.
@@
Set Samplers Two types of set samplers are adopted. The first type, named Walk-based Sampler, is to sample short-step random walks and eliminate repeated nodes during sampling. The second type, named Metric-based Sampler, is based on more principled graph metrics that measure the proximity between neighboring nodes and the seed node, such as personalized PageRank (PPR) scores (Jeh and Widom, 2003) or short path distances. Specifically, the walk-based sampler runs $M$-many $m$-step random walks, starting from each seed $u$ in parallel on the graph $\mathcal{G}$, and then puts only distinct nodes in sampled walks into set $\mathcal{S}_{u}$. The metric-based sampler, taking PPR-based (Bojchevski et al., 2020) as an example, first runs the push-flow algorithm (Andersen et al., 2006) to obtain an approximation of the PPR vector for each seed $u$, and then selects the top-$K$ nodes with the highest PPR scores into set $\mathcal{S}_{u}$. Mathematically, PPR scores are convergent landing probabilities of seeded random walks that reach infinite steps. Therefore, these two samplers complement each other by leveraging either more local or global structures of the graph. We use hyper-parameters $M$, $m$ to control random walks, and $K$ to control metric-based samplers, which are all set as some constants in practice. The complexity of the above offline sampling procedures is $O(|\mathcal{V}|)$.
@@
Structure Encoders The structure encoder is to construct structural features $\mathcal{Z}_{u,x}\in\mathbb{R}^{k}$ for each node $x$ in the sampled node set $\mathcal{S}_{u}$. These features prove to be crucial for inference tasks on graphs involving multiple nodes (Zhang et al., 2021) and can be conceptually understood as anchoring sampled node $x$ in the seed $u$’s neighborhood. One possible choice is landing probabilities of random walk (Li et al., 2019, 2020; Yin et al., 2022): each element $\mathcal{Z}_{u,x}[i]$ stores the counts of node $x$ landed at step $i$ of all walks rooted at the seed node $u$ divided by the total number of walks performed by the sampler. Landing probabilities (LPs) can be computed along with walk sampling. Another option is the shortest path distance (SPD) between $x$ and $u$ (Zhang and Chen, 2018; Li et al., 2020; Zhang et al., 2021), which records their relative distance in terms of reachability. PPR scores (Jeh and Widom, 2003) is another popular structural feature and can be naturally obtained by running a PPR-based sampler. Later, we denote the collection of structural features for all nodes in $\mathcal{S}_{u}$ as $\mathcal{Z}_{u}=\{\mathcal{Z}_{u,x}|x\in\mathcal{S}_{u}\}$.
@@
Figure 4. Node set $\mathcal{S}_{u}$ and its associated structural features $\mathcal{Z}_{u}$ stored in SpG. It contains three arrays to store the node set size, its elements, and associated structural features (with optional two-level indexing). Here, $D_{\text{SF}}$ shows the landing counts of nodes at different steps in sampled walks as an example, which can be normalized later to landing probabilities as structural features.
@@ 3.2. Set-based Storage - SpG
Node-set-based representations have advantages in terms of flexibility and compactness. However, the uneven sizes of sampled node sets pose great challenges to their storage and fast access. Note that these node sets are frequently visited in subsequent online phases for inference. To overcome these obstacles, SUREL+ designs a customized compressed sparse row (CSR) format called SpG, which reorganizes the storage of node sets and their structural features in a memory-efficient manner, as depicted in Fig. 4. Specifically, the node set $\mathcal{S}_{u}$ and its structural features $\mathcal{Z}_{u}$ are stored as a row of SpG, denoted as $\texttt{SpG}[u,:]$. Multiple node sets and their associated structural features are consolidated into three contiguous arrays:
@@
  • indices $I\in\mathbb{Z}^{\delta_{n+1}}$, a coalesce array of all node sets $\mathcal{S}_{u},~{}\forall u\in\mathcal{V}$. The segment $I[\delta_{u}:\delta_{u+1}]$ corresponds to the indices of sampled nodes of $\mathcal{S}_{u}$ stored in sorted order. This ordering is particularly useful for speeding up the join operation discussed in Sec. 3.3.
  • SFptr $D\in\mathbb{R}^{\delta_{n+1}}$, a pointer array contains the indices of the structural features $\mathcal{Z}_{u}$ stored in the array $D_{\text{SF}}$. The purpose of $D_{\text{SF}}$ is to eliminate duplicate structural features that typically reside in GPU memory. This secondary index can further reduce memory needs, especially when using LPs/SPDs that are likely to have many repeated values, but it is not necessary when using PPR scores since they tend to have distinct values.
@@
Regarding the cost of SpG, indptr array is of size $|\mathcal{V}|+1$, and the size of both indices and SFptr arrays is $\delta_{n+1}$. The compressed encoding array $D_{\text{SF}}$, has a size of $c*k$, where $c$ is the number of distinct structural features and $k$ denotes feature dimension. The overall complexity of this data structure is $O(|\mathcal{V}|+\delta_{n+1}+c*k)$.
Comparison with Other Methods Table 1 summarizes the space and time complexity comparison of GRL methods. By adopting the walk-based sampler (sampling $M$-many $m$-step walks), $\delta_{n+1}$ amounts to around one-fifth space used by SUREL of $O(mM|\mathcal{V}|)$, while the metric-based sampler (sampling top-$K$ PPR scores) results in $\delta_{n+1}=K|\mathcal{V}|$ and $K<mM$ in general. Both values are substantially lower than $O(S|\mathcal{E}|)$ used by SEAL, where $S$ is the average size of sampled subgraphs. For hosting structural features, SUREL+ employs the secondary index SFptr $D$ and retains only distinct structural features in $D_{\text{SF}}$ to further reduce the memory footprint. In practice, $c$ typically remains independent of $|\mathcal{V}|$. Equipped with SpG, SUREL+ is capable of handling industry-level graph data.
@@ 3.3. Joining Node Sets via Sparse Operations
The goal of joining node sets is to connect queried nodes and construct query-induced subgraphs from pre-sampled node sets to make predictions for SGRL queries. For a given query $Q$, we merge relevant node sets $S_{u},\forall u\in Q$ into $\mathcal{S}_{Q}=\bigcup_{u\in Q}\mathcal{S}_{u}$ and join their node-level structural features $\mathcal{Z}_{u}$ to the query level. In essence, query-level structural features $\mathcal{Z}_{Q}$ serve as an anchor that indicates the position of node $x\in\mathcal{S}_{Q}$ within the query-induced subgraph $\mathcal{G}_{Q}$. Specifically, for a node $x$ in $\mathcal{S}_{Q}$, the query-level structural features $\mathcal{Z}_{Q,x}$ are obtained by joining its node-level structural features $\mathcal{Z}_{u,x}$ for all queried node $u$ in $Q$ as
@@
where ||||| | denotes concatenation. In cases where $\mathcal{Z}_{u,x}$ does not exist as node $x\notin S_{u}$ while $u\in Q$, it is set to all zeros. For instance, in Fig. 5, node $b$ is in $\mathcal{S}_{v}$ but not in $\mathcal{S}_{u}$, hence $\mathcal{Z}_{u,b}$ is set to zero. $\mathcal{Z}_{Q}$ is a collection of $\mathcal{Z}_{Q,x},\forall x\in\mathcal{S}_{Q}$. Together, $\mathcal{S}_{Q}$ and $\mathcal{Z}_{Q}$ function as the query-induced subgraph $\mathcal{G}_{Q}$, which is fed into the neural encoder later to obtain embeddings for prediction.
The JOIN operator in databases is used to merge tables and establish connections. Concatenation in Eq. (1) requires matching among different node sets with varying sizes and arbitrary node orders, where an outer JOIN is well-suited for this task. In this case, the JOIN operator returns associated values based on the node index (specified common field) from target sets, regardless of its existence. To obtain $\mathcal{Z}_{Q,x}$, node sets $\{\mathcal{S}_{u}\}_{u\in Q}$ are treated as tables: if the index of a node $x$ matches one of the node indices in $\mathcal{S}_{u}$ for all $u$ in $Q$, then the associated structural features $\mathcal{Z}_{u,x}$ are appended; otherwise, the field is filled with zeros. However, iterating over all $\mathcal{S}_{u}$’s to retrieve $\mathcal{Z}_{u,x}$ for each node $x\in\mathcal{S}_{Q}$ is highly inefficient, as its complexity can be $O(|Q|*|\mathcal{S}_{Q}|^{2})$ per query. This becomes even more challenging when performing these operations for massive queries with varying sizes of $\mathcal{S}_{u}$ and $\mathcal{S}_{Q}$.
To tackle this issue, we design an efficient arithmetic operator SpJoin to perform joins on sparse data objects of SpG in parallel. This operator reduces the time complexity to $O(|Q|*|\mathcal{S}_{Q}|)$ per query by taking advantage of node indices of $\mathcal{S}_{u}$ stored in SpG are unique and in sorted order. The following demonstrates the use of SpJoin for a query $Q=\{u,v\}$.
Sparse Join Operator The operator SpJoin performs an outer JOIN on the sampled node sets from seeds $u$ and $v$ of query $Q$ stored in SpG as $\texttt{SpG}[u,:]$ and $\texttt{SpG}[v,:]$ through
@@
${\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[u,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1}~{}||~{}{\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[v,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1}$
@@
    It utilizes sparse arithmetic operations from SciPy (Virtanen et al., 2020): SpAdd performs an element-wise addition ($X\oplus Y$) of the non-zero elements in $X$ and $Y$; the resulting values are converted to binary via the bool operator and saved in the mask, which corresponds to node indices of the set $\mathcal{S}_{Q}$.
@@
    SpAdd are applied between mask and each $\texttt{SpG}[u,:],\,\forall u\in Q$ following by the reduction ‘-1’, which explicitly adds missing values (all zeros at default) to structural features $\mathcal{Z}_{u,x}$ for all $x$ if $x\not\in\mathcal{S}_{u}$ while $x\in\mathcal{S}_{v}$, for some $v\in Q\backslash\{u\}$.
@@
    When the secondary index SFptr $D$ is used, the results of SpJoin are pointers saved in SFptr, which can be used to gather the values of structural features $\mathcal{Z}_{Q}$ from the array $D_{\text{SF}}$.
Multithreading is employed to leverage the single program multiple data pattern in sparse operations of SpJoin. Since the processing time of each query linearly depends on the size of $\mathcal{S}_{Q}$, we further divide queries of each training batch into groups with nearly balanced sums of $|\mathcal{S}_{Q}|$’s, and assign one thread process a group to mitigate potential delays caused by uneven workloads.
@@
Comparison with SUREL (Yin et al., 2022) SUREL adopts hash-based join operators to build query-level structural features, but its overall computation and memory cost is much higher compared to SUREL+. This is due to the presence of numerous repeated nodes in walks (over half as Fig. 3 (b) shown). The set-based input of SUREL+ substantially reduces the workload of transferring data from CPU to GPU and also requires fewer per-node operations on GPU to process transmitted $\mathcal{Z}_{Q}$’s. As Table 1 shows, SUREL+ reduces time complexity from $O(mM)$ to $O(K)$, where $K<mM$ is the average size of sampled node sets. These advantages ultimately enable SUREL+ to achieve superior efficiency and scalability.

### 27 Dec 2023 v3

@@ 3. The framework of SUREL+
This section introduces SUREL+, whose key concept is to sample node sets and encode structural features offline and then join them online as a proxy of query-induced subgraph for representation learning. This approach only keeps distinct nodes in the sampled set for reuse in different queries, effectively addressing memory and computation concerns of node duplication in the walk-based representation adopted by SUREL (Yin et al., 2022). SUREL+ features a modular design that supports various set samplers, structure encoders, and set neural encoders to provide a trade-off between complexity and expressiveness after reducing subgraphs to node sets. Furthermore, SUREL+ introduces a customized sparse data structure SpG and an arithmetic operator SpJoin to store node sets and perform their online joins efficiently. Fig. 2 summarizes and compares SUREL+ and current SGRL models. The following subsections describe these modules in detail.
@@ 3.1. Set Samplers and Structure Encoders
SUREL+ uses set samplers to sample a set of nodes from each node’s neighborhood and calls structure encoders to construct the corresponding structural features. Both operations are executed offline: the former is primarily for computational benefits, while the latter is to offset the structure loss of node sets reducing from subgraphs (adopted by SEAL (Zhang and Chen, 2018; Zhang et al., 2021)) or walks (adopted by SUREL (Yin et al., 2022)). Conceptually, SUREL+ represents the node-induced subgraph $\mathcal{G}_{u}$ via a combination of (1) a node set $\mathcal{S}_{u}$ comprising unique nodes sampled from the neighborhood of node $u$ and (2) the associated structural features $\mathcal{Z}_{u}$ reflects the position of sampled nodes in $\mathcal{G}_{u}$.
@@
Set Samplers Two types of set samplers are adopted. The first type, named Walk-based Sampler, is to sample short-step random walks and then reduce sampled walks into a set of unique nodes. The second type, named Metric-based Sampler, is based on graph metrics that measure the proximity between neighboring nodes and the seed, such as personalized PageRank (PPR) scores (Jeh and Widom, 2003) or short path distances. Specifically, the walk-based sampler runs $M$-many $m$-step random walks, starting from each seed $u$ in parallel on the graph $\mathcal{G}$, and then only puts distinct nodes of sampled walks into the set $\mathcal{S}_{u}$. The metric-based sampler, taking PPR-based (Bojchevski et al., 2020) as an example, first runs the push-flow algorithm (Andersen et al., 2006) to obtain an approximation of the PPR vector for each seed $u$, and then selects the top-$K$ nodes with the highest PPR scores into the set $\mathcal{S}_{u}$. Mathematically, PPR scores are convergent landing probabilities of seeded random walks that reach infinite steps. Therefore, these two samplers complement each other by leveraging either more local or global structures of the graph. We use hyper-parameters $M$, $m$ to control random walks, and $K$ to control metric-based samplers, which are all set as some constants in practice. The complexity of the above offline sampling procedures is $O(|\mathcal{V}|)$.
@@
Structure Encoders The structure encoder is to construct structural features $\mathcal{Z}_{u,x}\in\mathbb{R}^{k}$ for each node $x$ in the sampled node set $\mathcal{S}_{u}$. These features prove to be crucial for inference tasks on graphs involving multiple nodes (Zhang et al., 2021) and can be conceptually understood as anchoring sampled node $x$ in the seed $u$’s neighborhood. One possible choice is landing probabilities of random walk (Li et al., 2019, 2020; Yin et al., 2022): each element $\mathcal{Z}_{u,x}[i]$ stores the counts of node $x$ landed at step $i$ of all walks rooted at the seed $u$ divided by the number of walks performed by the sampler. Landing probabilities (LPs) can be computed along with walk sampling. Another option is the shortest path distance (SPD) between $x$ and $u$ (Zhang and Chen, 2018; Li et al., 2020; Zhang et al., 2021), which records their relative distance in terms of reachability. PPR scores (Jeh and Widom, 2003) is also a popular structural feature and can be naturally obtained by running a PPR-based sampler. Later, we denote the collection of structural features for all nodes in $\mathcal{S}_{u}$ as $\mathcal{Z}_{u}=\{\mathcal{Z}_{u,x}|x\in\mathcal{S}_{u}\}$.
@@
Figure 4. Node set $\mathcal{S}_{u}$ and its associated structural features $\mathcal{Z}_{u}$ stored in SpG. It contains three arrays to store the size of node sets, sampled node indices, and associated structural features (with optional two-level indexing). Here, $D_{\text{SF}}$ shows the landing counts of nodes at different steps in sampled walks as an example, which can be normalized later to landing probabilities as structural features.
@@ 3.2. Set-based Storage - SpG
Set-based subgraph representation has advantages in terms of flexibility and compactness. However, the uneven sizes of sampled node sets pose great challenges to their storage and fast access. Note that, these node sets need to be frequently visited in subsequent online phases. To overcome these obstacles, SUREL+ designs a customized compressed sparse row (CSR) format called SpG, which reorganizes the storage of node sets and their structural features in a memory-efficient manner, as depicted in Fig. 4. Specifically, the node set $\mathcal{S}_{u}$ and its structural features $\mathcal{Z}_{u}$ are stored as a row of SpG, denoted as $\texttt{SpG}[u,:]$. Multiple node sets and their associated structural features are consolidated into three contiguous arrays:
@@
  • indices $I\in\mathbb{Z}^{\delta_{n+1}}$, a coalesce array of all node sets $\mathcal{S}_{u},~{}\forall u\in\mathcal{V}$. The segment $I[\delta_{u}:\delta_{u+1}]$ corresponds to node indices of the set $\mathcal{S}_{u}$ in sorted order. This ordering is particularly useful for speeding up the join operation discussed in Sec. 3.3.
  • SFptr $D\in\mathbb{R}^{\delta_{n+1}}$, an array contains the values of the structural features $\mathcal{Z}_{u}$, or the indices of encoding stored in the array $D_{\text{SF}}$. $D_{\text{SF}}$ is introduced to eliminate duplicate encoding of structural features, which typically reside in GPU memory. This two-level indexing can further reduce memory needs when LPs/SPDs are used, as they are likely to have many repeated values, but it is not necessary when using PPR scores since their values tend to be distinct.
@@
Regarding the cost of SpG, indptr array is of size $|\mathcal{V}|+1$, and the size of both indices and SFptr arrays is $\delta_{n+1}$. The compressed encoding array $D_{\text{SF}}$ has a size of $c*k$, where $c$ is the number of distinct structural features and $k$ denotes feature dimension. The overall complexity of SpG is $O(|\mathcal{V}|+\delta_{n+1}+c*k)$.
Comparison with Other Methods Table 1 summarizes the space and time complexity of GRL methods. By adopting the walk-based sampler (sampling $M$-many $m$-step walks), $\delta_{n+1}$ amounts to around one-fifth space of $O(mM|\mathcal{V}|)$ used by SUREL. The metric-based sampler (sampling top-$K$ PPR scores) results in $\delta_{n+1}=K|\mathcal{V}|$ and $K<mM$ in general. Both values are substantially lower than $O(S|\mathcal{E}|)$ used by SEAL, where $S$ is the average size of extracted subgraphs. SUREL+ further reduces the memory footprint, when the two-level indexing is employed for hosting structural features and only distinct values are stored in $D_{\text{SF}}$. In practice, $c$ typically remains independent of $|\mathcal{V}|$. SpG enables SUREL+ to handle SGRL tasks more efficiently on large-scale graph data.
@@ 3.3. Joining Node Sets via Sparse Operations
The goal of joining node sets is to connect queried nodes and construct query-level subgraphs from pre-sampled node sets to make predictions. For a given query $Q$, we merge relevant node sets $S_{u},\forall u\in Q$ into $\mathcal{S}_{Q}=\bigcup_{u\in Q}\mathcal{S}_{u}$ and join their node-level structural features $\mathcal{Z}_{u}$ to the query level. In essence, query-level structural features $\mathcal{Z}_{Q}$ record the relative position of each node $x\in\mathcal{S}_{Q}$ with respect to the set of queried nodes $Q$ (equivalently labeling the query-induced subgraph $\mathcal{G}_{Q}$). Specifically, for a node $x$ in $\mathcal{S}_{Q}$, the query-level structural feature $\mathcal{Z}_{Q,x}$ is obtained by merging its node-level ones $\mathcal{Z}_{u,x}$ for all queried node $u$ in $Q$ as
@@
where ||||| | denotes concatenation. In cases where $\mathcal{Z}_{u,x}$ does not exist as node $x\notin S_{u}$, it is set to all zeros. For instance, in Fig. 5, node $b$ is in $\mathcal{S}_{v}$ but not in $\mathcal{S}_{u}$, hence $\mathcal{Z}_{u,b}$ is set to zero. $\mathcal{Z}_{Q}$ is a collection of $\mathcal{Z}_{Q,x},\forall x\in\mathcal{S}_{Q}$. Together, $\mathcal{S}_{Q}$ and $\mathcal{Z}_{Q}$ function as the query-induced subgraph $\mathcal{G}_{Q}$, which is later fed into the neural encoder to obtain the query-level readout for prediction.
The JOIN operator in databases is used to merge tables and establish connections. Concatenation in Eq. (1) requires matching among different node sets with varying sizes and arbitrary node orders, where an outer JOIN is well-suited for this task. In this case, the JOIN operator returns associated values from target sets based on node indices as the specified common field, regardless of their existence. To obtain $\mathcal{Z}_{Q,x}$, node sets $\{\mathcal{S}_{u}\}_{u\in Q}$ are treated as tables: if the index of node $x$ matches one of the node indices in $\mathcal{S}_{u}$ for all $u$ in $Q$, then the associated structural feature $\mathcal{Z}_{u,x}$ is appended; otherwise, the field is filled with zeros. However, iterating over all $\mathcal{S}_{u}$’s to retrieve $\mathcal{Z}_{u,x}$ for each node $x\in\mathcal{S}_{Q}$ is highly inefficient, as its complexity can be $O(|Q|*|\mathcal{S}_{Q}|^{2})$ per query. This becomes even more challenging when performing these operations for massive queries with varying sizes of $\mathcal{S}_{u}$ and $\mathcal{S}_{Q}$.
To tackle this issue, we design an efficient arithmetic operator SpJoin to perform joins on sparse data objects of SpG in parallel. This operator reduces the per-query time complexity to $O(|Q|*|\mathcal{S}_{Q}|)$ by taking advantage of node indices of $\mathcal{S}_{u}$ stored in SpG are unique and in sorted order. The following demonstrates the use of SpJoin for a query $Q=\{u,v\}$.
Sparse Join Operator The operator SpJoin performs an outer JOIN for query $Q$ on the sampled node sets from seeds $u$ and $v$ stored in SpG as $\texttt{SpG}[u,:]$ and $\texttt{SpG}[v,:]$ through
@@
${\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[u,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1}~{}||~{}{\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[v,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1},$
@@
    It utilizes sparse arithmetic operations from SciPy (Virtanen et al., 2020): SpAdd performs an element-wise addition ($X\oplus Y$) of the non-zero elements in $X$ and $Y$; the resulting values are converted to binary via the bool operator and saved in the mask, which corresponds to node indices of the union set $\mathcal{S}_{Q}$.
@@
    SpAdd are applied between mask and each $\texttt{SpG}[u,:],\,\forall u\in Q$ following by the reduction ‘-1’, which explicitly adds missing values (all zeros by default) to structural features $\mathcal{Z}_{u,x}$ for all $x$ if $x\not\in\mathcal{S}_{u}$ while $x\in\mathcal{S}_{Q}$.
@@
    When the two-level indexing is enabled, the results of SpJoin are pointers saved in SFptr, which can be used to gather the values of structural features $\mathcal{Z}_{Q}$ from the array $D_{\text{SF}}$.
Multithreading is employed to leverage the pattern of single program multiple data in arithmetic operations of SpJoin. Since the processing time of each query linearly depends on the size of $\mathcal{S}_{Q}$, we further divide queries of each training batch into groups with nearly balanced sums of $|\mathcal{S}_{Q}|$’s, and assign one thread per group to mitigate potential delays caused by uneven workloads.
@@
Comparison with SUREL (Yin et al., 2022) SUREL adopts a hash-based join operator to construct query-level structural features, but its overall computation and memory cost is much higher than SUREL+. This is due to the presence of numerous repeated nodes in walks, depicted in Fig. 3 (b). The set-based input of SUREL+ substantially reduces the workload of transferring data from CPU to GPU and also requires fewer per-query operations on GPU to process transmitted $\mathcal{Z}_{Q}$’s. As Table 1 shows, SUREL+ reduces time complexity from $O(mM)$ to $O(K)$, where $K<mM$ is the average size of sampled node sets. These advantages ultimately enable SUREL+ to achieve superior performance in terms of efficiency and scalability.